# One protein is all you need

### v2

@@ G.3 Runtime performance
Figure A11: Hyperparameter search for protein structure prediction with ESM3 + ProteinTTT. We conducted a comprehensive grid search based on three key hyperparameters: learning rate (denoted as “lr”), number of gradient accumulation steps (denoted as “grad_steps”; with the batch size of two), and masking strategy (denoted as “mask”). We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the “beta30” and “betalinear30” distributions proposed in the ESM3 paper (Hayes et al., 2024). Each row in the table presents the mean TM-score and LDDT metrics with standard deviations across five random seeds on the CAMEO validation fold. The last row, denoted as “No ProteinTTT”, shows the performance of ESM3 without customization. The results indicate that ESM3 + ProteinTTT is robust to the choice of hyperparameters and consistently outperforms the base model across all configurations. We selected the configuration from the last row (excluding “No ProteinTTT”) to compute the results on the test fold. For the hyperparameter search, we used 30 customization steps instead of 50 to reduce computation time.

### v3

@@ Appendix I Limitations and future work
Figure A12: Magnitude of ESMFold pLDDT improvements after customization with ProteinTTT. The evaluation is performed on 317,882 proteins from the Big Fantastic Virus Database (BFVD). Percentage annotations indicate the fraction of proteins whose pLDDT increases by at least the corresponding value (e.g., 49.8% of proteins show an improvement of at least 5 pLDDT points).
Figure A13: Detailed analysis of ProteinTTT failure cases on the CAMEO test set. The figure shows the two entries that consistently exhibit a decrease in TM-score after customization with ProteinTTT across most random seeds (see Figure˜A7). (a) For chain B of PDB entry 7QII (white), the ground-truth structure is part of a dimer in which the conformation of chain B depends on interactions with chain A (black). In the monomeric prediction setting, this context is absent, making the precise helix arrangement inherently ambiguous. Both ESMFold and ESMFold + ProteinTTT correctly capture the helical composition but differ in the global configuration, leading to different TM-scores. (b) For chain A of PDB entry 7MQ4 (white), the reference structure is an NMR ensemble with substantial conformational variability (black). Both ESMFold and ESMFold + ProteinTTT recover the stable substructure (right part of the structure in black consisting of a helix surrounded by beta strands), yet produce different conformations in the flexible regions, where multiple arrangements are plausible.